OpenStudy (anonymous):

segment PR has an endpoint at (25, -5) and a midpoint of (18, -1). What is the value of the x-coordinate of the other endpoint?

OpenStudy (jdoe0001):

\(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ P&({\color{red}{ 25}}\quad ,&{\color{blue}{ -5}})\quad R&({\color{red}{ \square }}\quad ,&{\color{blue}{ \square }}) \end{array}\qquad \\ \quad \\ \left(\cfrac{{\color{red}{ \square }} + {\color{red}{ 25}}}{2}\quad ,\quad \cfrac{{\color{blue}{ \square }} + {\color{blue}{ (-5)}}}{2} \right)=(18,-1) \implies \cfrac{{\color{red}{ \square }} + {\color{red}{ 25}}}{2} = 18\)